Magic Company adds materials at the beginning of the process in Department A. The following information on physical units for Department A for the month of August is available.

Work in process, August 1 (60% complete with respect to conversion) 72,000

Started in August 480,000

Completed 480,000

Work in process, August 31 (30% complete with respect to conversion) 72,000

Required:

a. Compute the equivalent units for materials costs and for conversion costs using the weighted-average method.

b. Compute the equivalent units for materials costs and for conversion costs using the FIFO method.
Answer #1

a) Equivalent unit : Weighted average

Material Conversion
Unit transferred out 480000 480000
Ending work in process 72000 72000*30% = 21600
Equivalent unit 552000 501600

b) Equivalent unit : FIFO

Material Conversion
Beginning work in process 0 72000*40% = 28800
Unit Started and completed 408000 408000
Ending work in process 72000 72000*30% = 21600
Equivalent unit 480000 458400